Find all solutions of 24 = 4 + 4i Find all solutions of z2 = 3 - 2i
Answer #1

(1) Z=4+ui z = a + atbi = izl(cos(o)+ i since)) (Z1 = Ja²+br 2 = arbi Iz = N444 = 2 jy Os to tan (44) - z = UJI (cos( t) +at r=2 40 = 27712) ► DET I 2 = 1414 ( cosit tisin o) 72.=:-1,414 at rE 3 ; YO ILIT → tan Z z = 1.414 (cos 39+isin 37 ) Zz 3-l